OpenStudy (anonymous):

What is the missing number which makes the fractions equivalent? 3/5 = 15/? I will medal!

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ 3 }{ 5 }=\frac{ 15 }{ ? }\] Think... What times 3 = 15?

OpenStudy (anonymous):

Or \[\large \sf \frac{a}{b} = \frac{c}{d} \to ad=bc\] cuz cross multiplication

OpenStudy (anonymous):

Im confused. @LegendarySadist

pooja195 (pooja195):

3 x ? = 15 so basically 15/3=?

OpenStudy (anonymous):

15/3 = 5 @pooja195

pooja195 (pooja195):

Good so... \[\huge~\rm~\bf \frac{ 3 \times 5=15 }{ 5 \times 5 =? }\]

OpenStudy (anonymous):

25 @pooja195

pooja195 (pooja195):

so our fraction would be....?

OpenStudy (anonymous):

15/25? @pooja195

pooja195 (pooja195):

Perfect :)

OpenStudy (anonymous):

Thank you! @pooja195

pooja195 (pooja195):

yw ^_^